OpenStudy (anonymous):

if a1=0.4 &r=-2 find a15. if a1=-10 & d=0.8 find s20

OpenStudy (amistre64):

first one reads as geometric; the second reads as arithmetic

OpenStudy (anonymous):

which formula to use cause I got d instead of r on the first one

OpenStudy (amistre64):

r tends to mean ratio an=(a1)r^(n-1)

OpenStudy (amistre64):

a15 = .4 * (-2)^14

OpenStudy (amistre64):

s20 indicates the sum of all the numbers from a1 to a20 in the second case

OpenStudy (anonymous):

can you check my answer i got -48

OpenStudy (amistre64):

an=a1+d(n-1) is the arithmetic series

OpenStudy (amistre64):

-2 to an even power is positive

OpenStudy (anonymous):

so whats the formulae please

OpenStudy (amistre64):

youve asked 2 questions; which one are you working on?

OpenStudy (anonymous):

the first one at the moment

OpenStudy (amistre64):

ok, since r^0 = 1; and we start with n=1 we need to subtract it down to 0 \[a_n=a_1*r^{n-1}\]

OpenStudy (anonymous):

thanks and for the second please

OpenStudy (amistre64):

the second one is a bit trickier since they want the sum of the first 20 numbers

OpenStudy (amistre64):

d(0) = 0 so n=1 needs to subtract down to 0 \[a_n=a_1+d(n-1)\]

OpenStudy (amistre64):

the sum of it tho is: \[S_n=\frac{a_1(a_1+d(n-1))}{n}\]

OpenStudy (amistre64):

thats spoose to be /2

OpenStudy (anonymous):

going back to the first one I got a guffy answer can you check it for me

OpenStudy (amistre64):

first*last/2 \[S_n=\frac{a_1(a_1+d(n-1))}{2}\]

OpenStudy (anonymous):

3276.8 is that correct

OpenStudy (amistre64):

\[.4*(-2)^{14}=6553.6\]

OpenStudy (anonymous):

ok got it thanks

OpenStudy (amistre64):

a1=-10 & d=0.8 find s20 \[S_{20}=\frac{-10(-10+.8(19))}{2}\]

OpenStudy (amistre64):

i get -26 for the second one :)

OpenStudy (anonymous):

ok i will check it

OpenStudy (anonymous):

I still get -48